As a technician in a large pharmaceutical research firm, you need to produce 450. mL of a potassium dihydrogen phosphate buffer solution of pH = 7.05. The pKa of H2PO4− is 7.21. You have the following supplies: 2.00 L of 1.00 M KH2PO4 stock solution, 1.50 L of 1.00 M K2HPO4 stock solution, and a carboy of pure distilled H2O. How much 1.00 M KH2PO4 will you need to make this solution? (Assume additive volumes.)
Answer #1

pH = 7.05

pKa = 7.21

pH = pKa + log [conjugate base / acid]

7.05 = 7.21 + log [HPO42- / H2PO4-]

[HPO42- / H2PO4-] = 0.692

[HPO42- + H2PO4-] = 0.450

[HPO42- / H2PO4-] = 0.692

[0.450 - x / x ] = 0.692

0.450 - x = 0.692 x

x = 0.266

volume of KH2PO4 = 266 mL